Question 7 (1 point) What price today would you expect to pay for a stock with 13 percent required rate of return, a constant dividend growth of 4%, and an annual dividend of $2.50 which will be paid one year from today? A) $27.78 B) $30.28 C) $31.10 D) $28.88 Save
media%2F44e%2F44e59b54-3a82-4069-a903-a7
media%2F467%2F4679a7b9-9ce2-4746-a49f-55
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
Q7 A $27.78
P0 = D1/(ke-g)
Where P0 is the price
Ke is the required return 13.00%
g is the growth rate 4.00%
D1 is dividend at end of year 2.5
P0 = 2.50/(13% - 4%)
P0 = 2.50/(9%)
P0 = 27.78
Q8
b blend of equity and debt used by the firm
Q10 A 19.9% as explained below
Cash outflows           100,000.00
Cash inflows             30,000.00
PVf for 6 Years = 100,000/30000                   3.3333
PVF at 19%                   3.4098
PVF at 20%                   3.3255
Diff at 1 %                   0.0843
IRR = 20% + (3.3255-3.3333)/.0843
IRR = 20% + (-.0078)/.0843 19.91%
PVF 19% PVF20%
                                                                    1.00 0.8403 0.8333
                                                                    2.00 0.7062 0.6944
                                                                    3.00 0.5934 0.5787
                                                                    4.00 0.4987 0.4823
                                                                    5.00 0.4190 0.4019
                                                                    6.00 0.3521 0.3349
PVF for 6 Years 3.4098 3.3255